OpenStudy (anonymous):

Divide: Express in Scientific Notation 4.2x10^6/2.1x10^3

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \dfrac{ 4.2 \times 10^6}{2.1\times 10^3}\)

ganeshie8 (ganeshie8):

whats the value of 4.2/2.1 ?

OpenStudy (anonymous):

2.1

ganeshie8 (ganeshie8):

\(\large \dfrac{ 4.2 \times 10^6}{2.1\times 10^3} = \dfrac{ 2.1 \times 10^6}{10^3}\)

ganeshie8 (ganeshie8):

next use below exponent property : \(\large \dfrac{a^m}{a^n} = a^{m-n}\)

OpenStudy (anonymous):

\[2\times10^{2}\]?

ganeshie8 (ganeshie8):

\(\large \dfrac{ 4.2 \times 10^6}{2.1\times 10^3} = \dfrac{ 2.1 \times 10^6}{10^3} = 2.1 \times 10^{6-3}\)

OpenStudy (anonymous):

Okay so C and D are out.

OpenStudy (anonymous):

2x10^3 or 2.1x10^3?

ganeshie8 (ganeshie8):

\(\large \dfrac{ 4.2 \times 10^6}{2.1\times 10^3} = \dfrac{ 2.1 \times 10^6}{10^3} = 2.1 \times 10^{6-3} = 2.1\times 10^3\)

OpenStudy (anonymous):

I wanna go with 2.1

OpenStudy (anonymous):

Yayy.